OpenStudy (anonymous):

Solve: |2n - 3| = 5 A. {-1, 4} B. {1, 4} C. {-1, 1} D. {1, -4}

OpenStudy (nurali):

|2n - 3| = 5 2n-3=5 or 2n-3=-5 2n=8 or 2n=-2 n=4 or n=-1 (4,-1)

OpenStudy (anonymous):

Thanks :)

OpenStudy (nurali):

Anytime.